# Gravitational Waves Through Time: Scientific Significance, Detection Techniques, and Recent Breakthroughs

### v1

@@ Abstract
This paper provides a detailed review of gravitational waves. We begin with a thorough discussion regarding the history of gravitational waves, beginning even before Albert Einstein’s theory of general relativity, highlighting important developments and milestones in the field. We then discuss the scientific significance of gravitational wave detections such as the verification of general relativity and key properties of black holes/neutron stars. We extend our analysis into various detection techniques including interferometer-based detectors (LIGO, Virgo, GEO600), pulsar timing arrays, and proposed space-based detectors (LISA, DECIGO, BBO). Finally, we conclude our review with a brief examination of the captivating event GW190521.
@@ I Introduction
In Section II of this paper, we provide a brief history of gravitational waves, beginning even before Einstein’s theory of general relativity. Subsequently, in Section III, we delve into the scientific implications and significance of detecting gravitational waves. Section IV is dedicated to an exploration of various detectors and analytical techniques employed in the study of gravitational waves. In Section V, we shine a spotlight on a recent and particularly captivating gravitational wave event, denoted as GW190521.
@@ II A Brief History of Gravitational Waves
Figure 1: Snapshot from Cervantes-Cota 2016 [3] depicting a sketch of Weber’s cylindrical detector and a photo of Joseph Weber at the antenna.
@@
Figure 2: Snapshot from Weber 1969 [10] presents a table that lists some of the “coincidences” along with the period of time which must elapse for a coincidence with equal to or greater amplitudes to occur accidentally.
At the dawn of the 21st century, several state-of-the-art gravitational wave detectors, now integral to our scientific landscape, reached completion. For interferometer-based detectors, notable names include LIGO in the USA, Virgo in Italy, and GEO600 in Germany [3]. Other types of detectors such as pulsar timing arrays and space-based detectors have also become more prevalent in recent years and provide yet another avenue for gravitational wave detection. In Section IV, we will delve into a comprehensive exploration of each individual detector. Before this, however, we must focus on the scientific significance inherent in the study of gravitational wave detections.
@@ III Scientific Significance of Gravitational Wave Detections
As we delve into the intricacies of gravitational wave research, this section strives to elucidate the diverse spectrum of discoveries that GW have ushered into our scientific understanding. From confirming the predictions of Einstein’s theory of general relativity to understanding black holes and neutron star mergers, the exploration of gravitational wave detections opens unprecedented avenues for humanity’s understanding of the cosmos.
@@ III.1.1 Residuals Test
[…]
@@ III.2.2 Neutron Stars
[…]
@@
[…]
Figure 4: Snapshot from Sieniawska and Bejger 2019 [15] depicting different gravitational wave detectors and their respective range of data collection. The pink outline represents the GW150914 signal strain’s amplitude ($h_{0}$) and frequency range, whereas the blue region represents the ranges for pulsars.
@@ IV.1 Interferometer-based detectors
Figure 5: Snapshot from Caltech/MIT/LIGO Lab [19] depicting a diagram of LIGO’s interferometers with a gravitational waves approaching the detector
@@ IV.3.1 LISA
Figure 6: Snapshot taken from NASA [27] depicting the various detection techniques used to probe the gravitational wave spectrum.
@@
Figure 7: Snapshot taken from Danzmann and Rudiger 2003 [28] illustrating the proposed LISA mission. In the figure, the the triangle arms are scaled by a factor of 10.
@@ IV.3.2 Other proposed space-based missions
Figure 8: Snapshot taken Cutler and Holz 2009 [31] portraying the preliminary design for BBO. DECIGO would follow an identical composition, except the “arms” would be 50 times shorter than BBO’s.
@@ V Recent Breakthrough: GW 190521
In this last subsection, we delve into the remarkable scientific breakthrough represented by the gravitational wave event GW190521. This groundbreaking discovery, made possible by advanced gravitational wave observatories, was detected on May 21, 2019 by both LIGO and Virgo. Most notably, GW190521 was caused by the massive merger of two black holes: one at $\sim$85 solar-masses and the other at $\sim$66 solar-masses. If true, this would be the most massive black hole merger observed by our detectors to date [32].

### v2

@@ Abstract
This article, produced as part of an undergraduate research class, aims to provide an overview of gravitational waves, though it does not offer a comprehensive review. We begin with a brief discussion regarding the history of gravitational waves, beginning even before Albert Einstein’s theory of general relativity, highlighting important developments and milestones in the field. We then discuss the scientific significance of gravitational wave detections such as the verification of general relativity and key properties of black holes/neutron stars. We extend our analysis into various detection techniques including interferometer-based detectors (LIGO, Virgo, GEO600), pulsar timing arrays, and proposed space-based detectors (LISA, DECIGO, BBO). Finally, we conclude our review with a brief examination of the captivating event GW190521.
@@ I Introduction
In Section II of this paper, we provide a brief history of gravitational waves, beginning even before Einstein’s theory of general relativity. Subsequently, in Section III, we dive into the scientific implications and significance of detecting gravitational waves. Section IV is dedicated to an exploration of various detectors and analytical techniques employed in the study of gravitational waves. In Section V, we shine a spotlight on a recent and particularly captivating gravitational wave event, denoted as GW190521.
@@ II A Brief History of Gravitational Waves
Figure 1: Image from Cervantes-Cota 2016 [3] depicting a sketch of Weber’s cylindrical detector and a photo of Joseph Weber at the antenna.
@@
Figure 2: Table from Weber 1969 [10] listing some of the “coincidences” along with the period of time which must elapse for a coincidence with equal to or greater amplitudes to occur accidentally.
At the dawn of the 21st century, several state-of-the-art gravitational wave detectors, now integral to our scientific landscape, reached completion. For interferometer-based detectors, notable names include LIGO in the USA, Virgo in Italy, and GEO600 in Germany [3]. Other types of detectors such as pulsar timing arrays and space-based detectors have also become more prevalent in recent years and provide yet another avenue for gravitational wave detection. In Section IV, we provide a comprehensive exploration of each individual detector. Before this, however, we must focus on the scientific significance inherent in the study of gravitational wave detections.
@@ III Scientific Significance of Gravitational Wave Detections
This section strives to elucidate the diverse spectrum of discoveries that GW have ushered into our scientific understanding. From confirming the predictions of Einstein’s theory of general relativity to understanding black holes and neutron star mergers, the exploration of gravitational wave detections opens unprecedented avenues for humanity’s understanding of the cosmos.
@@ III.1.1 Residuals Test
Equations (1) and (2) are given by [12]. Equation (1) represents a generic form to calculate the fitting factor with h as the model and s as the signal. Equation (2) is a more specific form of Equation (1) with “90” referring to the 90% chance that the Signal-to-Noise Ratio (SNR) of the leftover signal (residual) after removing the best-fit GR template is less than or equal to a specific value denoted as SNR90.
Figure 3: Plot from Abbott et al. 2021 [13] showing the SNR90 for each GW event, as a function of the SNR maximum-likelihood template (SNRGR), with the corresponding p-values.
In Figure 3, the set of p-values show no distinct correlation between SNR90 and SNRGR. This implies that the SNR90 magnitude is solely influenced by the instrumental properties associated with each event and not influenced by the subtracted template (general relativity). This lack of correlation provides remarkably clear evidence in favor of general relativity [13].
@@ III.2.2 Neutron Stars
Specifically, our primary emphasis will be on the asymmetrical deformations within the neutron star. Magdalena Sieniawska and Michał Bejger in 2019 provide a set of equations to estimate a GW strain amplitude h0 from such neutron stars. They reach a succinct result which is given by Equation (3) [15].
@@
In Equation (3), $\epsilon$ represents a dimensionless parameter that depends on the level of asymmetry for the neutron star. The moment of inertia along a rotational axis is I3 which is dependent on the neutron star’s mass M and radius R (I3 $\approx$ MR2). $\nu$ is the neutron star’s rotational frequency and d represents the distance to the source (in this case, the neutron star) [15].
Figure 4: Figure from Sieniawska and Bejger 2019 [15] depicting different gravitational wave detectors and their respective range of data collection. The pink outline represents the GW150914 signal strain’s amplitude ($h_{0}$) and frequency range, whereas the blue region represents the ranges for pulsars.
@@ IV.1 Interferometer-based detectors
Figure 5: Diagram from Caltech/MIT/LIGO Lab [19] depicting a diagram of LIGO’s interferometers with a gravitational waves approaching the detector
@@ IV.3.1 LISA
Figure 6: Image from NASA [27] depicting the various detection techniques used to probe the gravitational wave spectrum.
@@
Figure 7: Illustration from Danzmann and Rudiger 2003 [28] illustrating the proposed LISA mission. In the figure, the the triangle arms are scaled by a factor of 10.
@@ IV.3.2 Other proposed space-based missions
Figure 8: Illustration from Cutler and Holz 2009 [31] portraying the preliminary design for BBO. DECIGO would follow an identical composition, except the “arms” would be 50 times shorter than BBO’s.
@@ V Recent Breakthrough: GW 190521
In this last subsection, we dive into the remarkable scientific breakthrough represented by the gravitational wave event GW190521. This groundbreaking discovery, made possible by advanced gravitational wave observatories, was detected on May 21, 2019 by both LIGO and Virgo. Most notably, GW190521 was caused by the massive merger of two black holes: one at $\sim$85 solar-masses and the other at $\sim$66 solar-masses. If true, this would be the most massive black hole merger observed by our detectors to date [32].